Random variable X has the following probability function:

f(x) = x – x3/4    for 0 < x < 2

Find P(X=0.8).
Answer #1

Ans:

Probability density function:

f(x)=P(x)=x-(x^3/4)

P(x=0.8)=0.8-(0.8^3/4)

P(x=0.8)=0.672